### v1

@@ 3.2 Replicating of existing techniques for experiments
At least 30 approaches adopt IR methods for localizing software bugs due to their computational efficiency and lightweight nature [22]. IR-based techniques rely on the textual similarity between bug reports and source code as a proxy of suspiciousness, which is explainable. We thus used three different baseline techniques from IR. We selected BugLocator [11] as our initial baseline technique since it is the seminal work on IR-based bug localization. Then, we chose BLUiR [12], recognized as the first structured IR-based technique that integrates code structure and bug report structure in its analysis. As our third baseline, we selected BLIA [23], notable for its incorporation of meta-components such as stack traces and version control history, which has demonstrated better performance compared to other IR-based techniques.
Unlike the above IR-based techniques that rely on the textual similarity between code and bug reports, DL-based approaches have the potential to uncover complex, non-linear relationships between dependent and independent variables [26, 27, 28]. Thus, we also used Deep Neural Network (DNN) for bug localization in deep learning systems, adapting an existing work – DNNLOC [17]. DNNLOC is a seminal work on DL-based hybrid approach to bug localization, integrating both traditional and deep learning techniques.
Most of the recent techniques are based on these four primary approaches with incremental improvements [14, 13, 47, 19]. We chose these baseline methods for our study, which can be considered as a representative sample of the existing approaches for bug localization.
@@
Since the original authors’ replication packages were unavailable, we used the publicly available versions to replicate BugLocator [49], BLUiR [49], and DNNLOC [50]. We also carefully adapted the BLIA from the original replication package [23] to our datasets. Our replication package is provided for further experimental details [46].
@@ 3.3.2 Mean Average Precision
Precision@K measures the precision of each buggy source document’s occurrence within a ranked list. Average Precision@K (AP) computes the average precision for all buggy documents within the ranked list against a search query (a.k.a., bug report). Mean Average Precision (MAP) is the average AP@K value across all queries in a system.
@@ 4 Study Finding
[…]
Comparison of the localization performance: Table 4 compares the performance of our baseline techniques in bug localization between deep learning systems and non-deep learning systems (a.k.a traditional software systems). We used three different evaluation metrics – Top@k, MRR, and MAP, for our comparative analysis.
We see notable differences in performance between the two types of systems when evaluating BugLocator, BLUiR, BLIA, and DNNLOC methods. In particular, for BugLocator, the differences in MAP and MRR are 31.59% and 29.46%, respectively. On the other hand, for BLUiR, the differences in MAP and MRR are also substantial, 33.25% and 30.24%, respectively. For BLIA, the differences in MAP and MRR are 34.14% and 30.77%, respectively. Finally, for DNNLOC, the bug localization performance for both cases has improved, but the gap remains large as the difference in MAP and MRR are 31.43% and 26.25%, respectively. We calculated the performance difference using the PerformanceDiff metric from Wattanakriengkrai et al. [51]. Fig. 2 visualizes the MAP measures, and the differences are clearly visible. Overall, the results show that all four approaches perform lower when localizing bugs in deep learning systems, and the trend is consistent across all three metrics.
Figure 2: Performance comparison of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC) between Deep Learning Systems (DLSW) and Non-Deep Learning Systems) (NDLSW)
We also perform statistical tests to determine the significance of the performance gap between the two types of systems – deep learning systems and non-deep learning systems (Table 5). We took the Reciprocal Rank (RR) and Average Precision (AP) results of all the samples for each of the four approaches. Then, we performed Shapiro-Wilk normality test [52], which reported normal distribution for those metrics. Then, we used appropriate significance and effect size tests to compare the result values from the two types of systems. For the normal distribution, we used t-test as the parametric test [33]. In all significance tests, the p-values were less than the threshold (0.05) for each of the four approaches. Thus, the null hypothesis can be rejected for all comparisons. In other words, the performances of all baseline techniques significantly differ between the two types of systems.
Table 4: Performance of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC) in bug localization
@@
Table 5: Statistical tests for the performance gap of existing approaches between the deep learning systems (DLSW) and non-deep learning systems (NDLSW)
@@
Method Metric Sig. (p-val) Effect Size BugLocator RR 0.00008721** Medium (0.3834) AP 0.00004211** Medium (0.4023) BLUiR RR 0.00395* Medium (0.2709) AP 0.00339* Medium (0.3154) BLIA RR 0.00000734*** Large (0.5012) AP 0.00000819*** Large (0.4896) DNNLOC RR 0.00000345*** Large (0.5237) AP 0.00000389*** Large (0.5110) Sig.: Significance, p-val: p-value, ***: Large, **: Medium, *: Small
@@
While the significance of a result indicates how probable it is that it is due to chance, the effect size indicates the extent of the difference [53]. Hence, we performed the Cohen’s D effect size test [34], and our analysis found a medium to large effect size for all cases (Table 5). Thus, our results from effect size tests reinforce the above finding from significance tests. In other words, the existing techniques perform significantly poorly in localizing bugs from deep-learning software systems. Even though our findings above match natural intuition, we performed extensive experiments using four different baselines, which resulted in strong empirical evidence. Thus, not only our findings reinforce the existing understanding and belief about the challenges of the bugs in deep learning systems, but also they substantiate them with solid empirical evidence and demonstrate the performance gap of existing solutions in localizing the two categories of bugs.
Comparison among the categories of deep learning systems: Our dataset, Denchmark, consists of deep learning systems from various classes, including frameworks, libraries, and tools. We focus on examining if there are any differences in bug localization performance across these system classes. Thus, we employed four existing techniques (BugLocator, BLUiR, BLIA, DNNLOC) to evaluate their bug localization performance for each class.
Table 6: Example of a bug report from deep learning framework [54]
Framework Bug (Bug ID: 10224)
Title
Language model example cannot be run
Description
The language model example cannot be run without manually
creating a data folder. There are also inconsistencies between the
documentation and the code. Optional argument – data DATA
location of the data corpus does not appear in the code train.py
Detailed BR: https://github.com/apache/mxnet/issues/10224
A deep learning framework is a software platform that provides the environment for designing, training, and deploying deep learning models [55]. Examples include TensorFlow [^1] https://www.tensorflow.org/, PyTorch [^2] https://pytorch.org/, and Apache MXNet [^3] https://mxnet.apache.org/versions/1.9.1/. These frameworks come with pre-defined modules and functions and offer a structured way to implement deep learning architectures using high-level programming interfaces [56]. The example bug in Table 6 [54] is characterized by the inability to run a language model without manually creating a data folder. It represents a framework bug because it directly impacts the core functionalities of the framework, specifically the design and training of models. Frameworks are expected to provide seamless, user-friendly environments for developing deep learning models, and issues that hinder the ease of use, such as documentation inaccuracies and additional manual setup steps, are indicative of problems at the framework level.
Table 7: Example of a bug from deep learning library [39]
Library Bug (Bug ID: 313)
Title
A bug in GPT2Tokenizer
Description
GPT2Tokenizer fails to recover a sentence
\”BART is a seq2seq model.\”
with encoded ids of it.
The output sentence is \”BART is a seqseq model.\”.
It should be related to numbers’ processing.
URL: https://github.com/tanyuqian/texar-pytorch
/blob/master/examples/bart/gpt2_tokenizer_bug.py
Figure 3: Performance comparison of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC) for various deep learning project classes
Table 8: Example of a bug from deep learning tool [58]
Tool Bug (Bug ID: 5596)
Title
Tensorboard can not load all Hyperparameters keys
Description
Encountered an issue in TensorBoard where it couldn’t load
all hyperparameter keys when I used writer.add-hparams with
different hparam-dict parameters in multiple experiments. This
problem made it difficult to track and compare different
hyperparameter settings across these experiments, affecting the
overall functionality of TensorBoard. I provided code snippets and
responses to showcase the inconsistency in the display of
hyperparameters in TensorBoard’s interface. This issue
is significant as it hinders the effective use of TensorBoard
for experiment tracking and analysis.

### v2

@@ 3.2 Replicating of existing techniques for experiments
Over the past decade, Information Retrieval (IR)-based bug localization techniques have gained significant attention, and some of these techniques have been proven effective in localizing bugs from traditional software systems [11, 12, 14, 53, 23]. Our extensive literature review shows that more than 30 peer-reviewed research papers have utilized IR for bug localization and highlighted its benefits [22, 54]. These techniques rely on textual similarities between bug reports and source code to identify relevant files or methods. Our motivation was to explore whether the success of these well-established IR-based methods could extend to DL systems.
While spectrum-based bug localization techniques are widely studied and capable of method-level or line-level localization [55, 56], they might face significant challenges when applied to DL systems. Specturm-based techniques analyze program traces from passing and failing tests, but DL systems often present complexities such as non-deterministic behavior, silent bugs (bugs that do not cause crashes or generate stack traces) [36, 57], and external dependencies (e.g., GPUs), which make obtaining such traces difficult or computationally prohibitive. In contrast, IR-based techniques do not require execution traces and instead rely on textual similarity between bug reports and code, which can be particularly effective in DL projects where bugs often stem from code, configuration, or data script issues [58]. Although IR techniques localize bugs at a coarser file level, this granularity serves as a practical initial step for large-scale DL projects, narrowing down the search space in extensive codebases and complex systems with multiple components [59]. Given these advantages and the computational efficiency of IR-based approaches, we focused our study on evaluating their applicability in DL systems.
We selected three IR-based techniques as our baselines for bug localization: BugLocator [11], BLUiR [12], and BLIA [23]. These techniques represent critical advancements in the field. BugLocator is a seminal work that first introduced the application of Information Retrieval (IR) through the Vector Space Model (VSM) for bug localization. BugLocator paved the way for subsequent research in IR-based localization by using both source code and historical bug reports. On the other hand, BLUiR is built upon BugLocator by introducing structured information retrieval, which integrates structural elements from both source code and bug reports, marking a significant improvement over earlier methods reliant solely on textual similarity. Thirdly, BLIA extended BLUiR’s approach by incorporating additional meta-components such as stack traces, code change histories, and version control information. This combination has shown superior performance over earlier techniques like Locus [60], AmaLgam [14], and BRTracer [53], making BLIA a comprehensive and effective baseline for our study. These three techniques not only represent key advancements in IR-based bug localization but are also computationally efficient, making them feasible for large-scale DL projects where obtaining execution traces is often impractical.
In addition to IR-based techniques, we selected DNNLOC [17], a hybrid approach that combines traditional IR methods with Deep Neural Networks (DNNs). Unlike pure IR-based methods, DNNLOC leverages the ability of deep learning models to capture non-linear and complex relationships between bug reports and source code. This hybrid approach brings the computational efficiency of IR-based methods while enhancing localization through deep learning’s ability to uncover hidden patterns in large datasets. Moreover, we included bjXnet [30], a state-of-the-art model that applies an attention mechanism with a code property graph to localize buggy files. This ensures that our study covers a wide range of methodologies, from traditional IR-based approaches to more advanced deep learning methods with attention mechanisms.
We selected these five techniques to ensure a comprehensive evaluation of bug localization methodologies. They represent a spectrum of approaches, ranging from early IR-based methods to cutting-edge deep learning models, each contributing novel ideas to the field. We chose these baselines based on their relevance to our study goals, computational feasibility, representativeness of the existing literature, and the interpretability of their results. Even DNNLOC, despite incorporating a deep learning component, maintains computational efficiency due to its hybrid nature. The explainability of IR-based approaches offers another significant advantage, as developers can easily interpret the results based on textual similarities—a crucial factor when debugging complex DL systems. The details of each baseline are as follows.
@@
bjXnet [30] employs a multimodal representation learning framework that integrates the Code Property Graph (CPG) [61] and attention mechanism [62] to improve bug localization. CPG provides a rich, structured representation of the source code by incorporating different aspects such as abstract syntax trees (ASTs), control flow graphs (CFGs), and data flow graphs (DFGs). This comprehensive representation captures both the syntactic and semantic properties of the code, facilitating a deeper understanding that is crucial for identifying buggy parts. The attention mechanism in bjXnet allows the model to focus on the most critical components within the CPG, effectively highlighting important patterns and dependencies that are likely associated with bugs. By weighting specific nodes or edges more heavily, bjXnet can localize bugs more accurately than methods relying solely on textual similarity. bjXnet has outperformed several earlier techniques - such as BugLocator [11], DeepLocator [19], DeepLoc [63], and DreamLoc [64]. Given its cutting-edge methodology and empirical success, we selected bjXnet for our empirical study.
Since the original authors’ replication packages were unavailable, we used the publicly available versions to replicate BugLocator [11], BLUiR [12], and DNNLOC [17, 65] that have been validated in peer-reviewed papers [54]. We also carefully adapted the BLIA [23] and bjXnet [30] from their original replication packages to our datasets. Our replication package is provided for further experimental details [66].
We validated our replication of the baseline techniques by conducting experiments on a common subject system – SWT. All baseline techniques [11, 12, 23, 17, 30] utilized the SWT dataset to assess performance in their original work, which makes it a viable choice for our validation.
Model MAP (Original) MAP (Replication) Difference (%)
BugLocator 0.45 0.44 2.22
BLUiR 0.58 0.57 1.72
BLIA 0.65 0.63 3.08
DNNLoc 0.37 0.36 2.70
bjXnet 0.55 0.53 3.64
Table 7: Comparison of Original and Replicated Techniques in Terms of MAP Performance on the SWT Dataset
From Table 7, we notice that the MAP differences range from 1.72% to 3.64% and fall within the margin of error, confirming the reliability of our replications. We also conducted a paired t-test, which resulted in a p-value of 0.12, showing the differences are statistically insignificant and validating the accuracy of our replication process.
Our study focuses on localizing bugs in deep learning systems, which are predominantly Python-based. To maintain consistency and eliminate confounding factors in our comparative analysis, we also examine traditional software systems written in Python. Although the baseline techniques [11, 12, 23, 17, 30] were initially developed for Java projects, recent empirical evidence suggests that their core principles are largely language-agnostic [67]. However, none of these techniques rely on language-specific features for bug localization. Our primary objective is to evaluate the effectiveness of existing bug localization techniques in deep learning systems. The results we obtained from traditional Python projects (NDLSW) are comparable to those reported for Java projects when these techniques were first introduced (Table 8). This comparability suggests that the programming language may not significantly influence the validity or applicability of our findings. By focusing on Python, we address the predominant language in deep learning while ensuring our results remain relevant to a broader context of software engineering.
@@ 3.3.2 Mean Average Precision
Precision@K measures the precision of each buggy source document’s occurrence within a ranked list. Average Precision@K (AP) computes the average precision for all buggy documents within the ranked list against a search query (a.k.a. bug report). Mean Average Precision (MAP) is the average AP@K value across all queries in a system.
@@ 4 Study Finding
4.1 Answering RQ1: How effective are the existing approaches in localizing bugs from deep learning systems?
4.1.1 Performance comparison between DLSW and NDLSW
Table 8 compares the performance of our baseline techniques in bug localization between deep learning systems and non-deep learning systems (a.k.a traditional software systems). We used three different evaluation metrics – Top@k, MRR, and MAP, for our comparative analysis.
Our results reveal notable performance differences between the two types of systems across all five methods. Specifically, for BugLocator, the differences in MAP and MRR are 31.59% and 29.46%, respectively. BLUiR exhibits differences of 33.25% in MAP and 30.24% in MRR. BLIA shows even larger difference, with MAP and MRR differences of 34.14% and 30.77%, respectively. DNNLOC, while improving overall performance, still presents significant gaps with differences of 31.43% in MAP and 26.25% in MRR. Notably, bjXnet, despite being a state-of-the-art technique that leverages advanced features like Code Property Graphs (CPG) and attention mechanisms, also demonstrates a substantial performance gap. The differences in MAP and MRR for bjXnet are 27.87% and 24.24%, respectively, which is consistent with the performance gaps observed in other techniques included in our study.
Figure 2: Performance comparison of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) between Deep Learning Systems (DLSW) and Non-Deep Learning Systems) (NDLSW)
We calculated these performance differences using the PerformanceDiff metric from Wattanakriengkrai et al. [68]. Fig. 2 visualizes the MAP measures, and the differences are clearly visible. Overall, the results show that all five approaches perform lower when localizing bugs in deep learning systems, and the trend is consistent across all three metrics.
We also perform statistical tests to determine the significance of the performance gap between the two types of systems – deep learning systems and non-deep learning systems (Table 9). We took the Reciprocal Rank (RR) and Average Precision (AP) results of all the samples for each of the five approaches. Then, we performed Shapiro-Wilk normality test [69], which reported normal distribution for those metrics. Then, we used appropriate significance and effect size tests to compare the result values from the two types of systems. For the normal distribution, we used t-test as the parametric test [34]. In all significance tests, the p-values were less than the threshold (0.05) for each of the five approaches. Thus, the null hypothesis can be rejected for all comparisons. In other words, the performances of all baseline techniques significantly differ between the two types of systems.
Table 8: Performance of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) in bug localization
@@ 4.1.1 Performance comparison between DLSW and NDLSW
Table 9: Statistical tests for the performance gap of existing approaches between deep learning systems (DLSW) and non-deep learning systems (NDLSW)
@@
Method Metric Sig. (p-val) Effect Size BugLocator RR 0.00008721** Medium (0.3834) AP 0.00004211** Medium (0.4023) BLUiR RR 0.00395* Medium (0.2709) AP 0.00339* Medium (0.3154) BLIA RR 0.00000734*** Large (0.5012) AP 0.00000819*** Large (0.4896) DNNLOC RR 0.00000345*** Large (0.5237) AP 0.00000389*** Large (0.5110) bjXnet RR 0.00000250*** Large (0.5400) AP 0.00000275*** Large (0.5280) Sig.: Significance, p-val: p-value, RR: Reciprocal Rank, AP: Average Precision, ***: Large, **: Medium, *: Small
@@
While the significance of a result indicates how probable it is that it is due to chance, the effect size indicates the extent of the difference [70]. Hence, we performed the Cohen’s D effect size test [35], and our analysis found a medium to large effect size for all cases (Table 9). Thus, our results from effect size tests reinforce the above finding from significance tests. Even the state-of-the-art technique bjXnet shows the gap is statistically significant with a large effect size. This indicates that the disparity is not only statistically meaningful but also practically significant, underscoring the challenges of localizing bugs in DL systems.
To better understand the challenges of bug localization in deep learning systems, we performed a deeper analysis and selected bjXnet, a state-of-the-art technique. Although it has outperformed other baselines, its performance remains inadequate for deep learning systems. Our analysis identified several key factors contributing to bjXnet’s reduced effectiveness in deep learning systems. By examining a random subset of bugs (100 bugs) from the Denchmark dataset, we found that DL systems heavily rely on high-level abstractions provided by frameworks like TensorFlow and PyTorch.
Figure 3: A code snippet from our DLSW dataset showing error in activation function
These abstractions include predefined components such as neural network layers (e.g., Dense, Conv2D), activation functions (e.g., relu, sigmoid), and optimizers (e.g., Adam, SGD). They encapsulate complex mathematical operations—like matrix multiplications and gradient computations—into single library calls. This encapsulation means that essential computational logic is hidden within these library functions rather than being explicitly defined in the source code, which might lead to bjXnet’s low performance because the technique relies on analyzing source code to trace errors. In 87% of the cases we examined, core functionalities such as model training, gradient calculations, and loss computations were implemented using these high-level library calls. For example, Fig. 3 shows a bug that involves the incorrect selection of activation functions in a dense layer—using ’sigmoid’ in regression tasks.
Moreover, DL systems build computation graphs where each layer and operation represents a node. In frameworks like TensorFlow, the computation graph is typically static, defined before runtime. However, in newer versions (e.g., TensorFlow [^7] https://www.tensorflow.org/api 2+ with eager execution) and other frameworks like PyTorch, computation graphs can evolve dynamically. This dynamic behavior contrasts with traditional systems, making it difficult for static analysis tools like bjXnet to trace errors. These tools struggle to handle the dynamic flow of data and operations, especially when key logic is abstracted in high-level library functions. Other IR-based baselines like BLUiR also have this limitation where only static code analysis is inadequate for DL systems.
In short, even though our findings above match natural intuition, we performed extensive experiments using five different baselines, which resulted in strong empirical evidence. Thus, not only do our findings reinforce the existing understanding and belief about the challenges of the bugs in deep learning systems, but also they substantiate them with solid empirical evidence and demonstrate the performance gap of existing solutions in localizing the two categories of bugs.
Potential Confounding Factors: To ensure fairness in comparison and highlight potential confounding factors, such as code scale, code complexity, and fault representativeness, which could influence the effectiveness of bug localization techniques, we conducted additional analyses as follows.
1. Code Scale: We analyzed the Lines of Code (LOC) from both datasets to investigate whether code scale biases our results. We found that the average LOC per file is higher for NDLSW (325.32) compared to DLSW (280), and the median LOC per file is also higher for NDLSW (244) than for DLSW (210). Such differences in LOC might pose a confounding factor in our experiment. To mitigate this, we repeated a part of our experiment using source documents with similar LOC.
We defined a consistent LOC threshold using the median values from both datasets (NDLSW and DLSW). The median was chosen as a robust central measure to handle the skewed distribution, a standard practice in similar studies [71]. We averaged the medians (244 and 210 LOC, respectively) to get a common median of 227 LOC, then applied a 25% margin (upper bound: 284 LOC, lower bound: 170 LOC) to define our LOC range. This thresholding method follows established guidelines for managing skewed data distributions [71]. We selected 439 files from NDLSW and 618 files from DLSW within the defined LOC threshold. Then, we randomly sampled 350 files from each dataset (95% confidence interval, 5% margin of error) to measure bug localization performance while controlling for code scale.
Table 10: Performance of bug localization approaches on the subsample of 350 files
Method Top@1 Top@5 Top@10 MRR MAP DLSW Subset BugLocator 0.325 0.523 0.600 0.355 0.300 BLUiR 0.180 0.450 0.550 0.290 0.240 BLIA 0.390 0.590 0.695 0.410 0.340 DNNLOC 0.450 0.670 0.770 0.440 0.390 bjXnet 0.470 0.690 0.780 0.460 0.410 NDLSW Subset BugLocator 0.400 0.650 0.770 0.500 0.450 BLUiR 0.285 0.550 0.670 0.420 0.350 BLIA 0.470 0.700 0.800 0.570 0.510 DNNLOC 0.580 0.750 0.830 0.580 0.560 bjXnet 0.600 0.770 0.850 0.600 0.580
Table 11: Statistical tests for the performance gap of existing approaches between DLSW subset and NDLSW subset
Method Metric Sig. (p-val) Effect Size BugLocator RR 0.00009123** Medium (0.3652) BugLocator AP 0.00005467** Medium (0.3824) BLUiR RR 0.00412* Small (0.2551) BLUiR AP 0.00374* Small (0.2992) BLIA RR 0.00000912*** Large (0.4850) BLIA AP 0.00000899*** Large (0.4723) DNNLOC RR 0.00000381*** Large (0.5101) DNNLOC AP 0.00000419*** Large (0.4956) bjXnet RR 0.00000280*** Large (0.5250) bjXnet AP 0.00000290*** Large (0.5120) Sig.: Significance, p-val: p-value, RR: Reciprocal Rank, AP: Average Precision, ***: Large, **: Medium, *: Small
We evaluated the performance of our bug localization techniques on these selected subsets. The results are presented in Table 10. The performance trends in the deep learning subset are similar to those observed in the full deep learning dataset. DNNLOC and bjXnet perform best across all metrics, followed by BLIA, BugLocator, and BLUiR. Although the overall performance is slightly lower due to the smaller sample size, our overall findings remain consistent. In the non-deep learning subset, all methods achieve higher scores compared to the deep learning subset, with DNNLOC and bjXnet performing significantly better. We also conducted statistical tests (refer to Table 11) to determine if the performance differences between DLSW and NDLSW are significant when the code scale confounding factor is mitigated. The p-values indicate that all methods’ performance differences between DLSW and NDLSW are statistically significant, with medium to large effect sizes.
To summarize, we controlled for differences in LOC between the two datasets. After re-running our experiments using comparable file sizes, the performance trends remained consistent, and the results showed that the differences between DLSW and NDLSW are still statistically significant. This indicates that our findings are not influenced by code scale variations.
Table 12: Code Complexity Metrics for deep learning systems (DLSW) and non-deep learning systems (NDLSW)
Dataset Median LOC AVG LOC AVG CC AVG Halstead AVG MI BuGL (NDLSW) 244 325.32 8.72 1456.23 75.30 Denchmark (DLSW) 210 280 9.5 1482.30 68.44 LOC: Lines of Code, CC: Cyclomatic Complexity, MI: Maintainability Index
2. Code Complexity: We also considered code complexity as a potential confounding factor. To compare the code complexity between DLSW (Denchmark) and NDLSW (BuGL), we used established software metrics: Cyclomatic Complexity (CC) [72], Halstead Complexity [73], and the Maintainability Index (MI) [74].
Cyclomatic Complexity measures the number of linearly independent paths through a program’s source code, providing insight into control flow complexity. Halstead Complexity assesses cognitive complexity based on the number of operators and operands, reflecting the effort required to understand the code. The Maintainability Index evaluates code maintainability by considering factors like LOC, CC, and Halstead metrics.
Our findings, summarized in Table 12, reveal that the average Cyclomatic Complexity per file is slightly higher for DLSW (9.5) compared to NDLSW (8.72). Similarly, the average Halstead Complexity is comparable between the two datasets (DLSW: 1482.30, NDLSW: 1456.23), and the Maintainability Index is slightly higher for NDLSW (75.30) than DLSW (68.44). We conducted statistical tests (e.g., independent two-sample t-tests) based on the sample size and the distribution of the data for each of the complexity metrics.
The t-test results are as follows: Cyclomatic Complexity: t(698) = 1.45, p = 0.15; Halstead Complexity: t(698) = 0.78, p = 0.44; Maintainability Index: t(698) = -1.32, p = 0.19. These p-values are all above 0.05, indicating that the differences in code complexity metrics between DLSW and NDLSW are not statistically significant.
Our analysis shows that both BuGL and Denchmark exhibit similar levels of code complexity. While their structures differ—BuGL being more modular and Denchmark relying on computational tasks—the statistical analysis of Cyclomatic Complexity, Halstead Complexity, and Maintainability Index indicates that both datasets are comparable in terms of code complexity.
Table 13: API Bug Localization Performance (Subset of 100 Bugs) for deep learning systems (DLSW) and non-deep learning systems (NDLSW)
Method MRR MAP DLSW Subset BugLocator 0.365 0.430 BLUiR 0.410 0.270 BLIA 0.430 0.400 DNNLOC 0.510 0.480 bjXnet 0.530 0.500 NDLSW Subset BugLocator 0.470 0.520 BLUiR 0.460 0.350 BLIA 0.520 0.490 DNNLOC 0.590 0.580 bjXnet 0.610 0.600 MRR: Mean Reciprocal Rank, MAP: Mean Average Precision
Figure 4: A code snippet from the DLSW dataset showing a PyTorch API misuse leading to a memory leak
3. Fault Representation: We also considered fault representativeness in our extended analysis. To address this concern, we focused on API bugs, which are common in both DLSW and NDLSW. API bugs, such as incorrect interactions with external libraries or frameworks, are prevalent in both deep learning systems (e.g., TensorFlow, PyTorch) and traditional software systems (e.g., Matplotlib, Requests). We identified 195 API bugs in DLSW and 102 in NDLSW through keyword searches and manual analysis by two authors. We selected 100 API bugs from each dataset and evaluated the bug localization performance of our techniques. Despite mitigating the confounding factor by focusing on API bugs, bug localization remains more challenging in DL systems. From Fig. 4, in PyTorch, missing ‘optimizer.zero_grad()‘ causes a memory leak, leading to out-of-memory errors after prolonged training, making the bug harder to locate using the baselines.
Our comprehensive analysis addresses potential confounding factors such as code scale, code complexity, and fault representation. Even after controlling for these factors, our findings show that bug localization remains significantly more challenging in deep learning systems compared to traditional software systems. This suggests that, despite the presence of similar bug types (e.g., API misuse), bugs in deep learning systems often manifest in more complex and indirect ways. As a result, traditional methods may be insufficient for effectively localizing bugs in deep learning systems, highlighting the need for new approaches tailored to their unique characteristics.
4.1.2 Comparison among the categories of deep learning systems
The original Denchmark dataset comprises 136 projects and 2,365 bug reports, categorized into Frameworks, Tools, Libraries, Applications, Engines, Platforms, and Compilers [32]. Our focus centered on the core components of deep learning systems—Frameworks, Tools, and Libraries—as these are essential and represent the majority of relevant bug reports. We aimed to examine whether there are differences in bug localization performance across these system categories, and we classified the dataset categories as follows.
Filter 1: From the original dataset, we selected 25 Frameworks, 31 Tools, and 44 Libraries, totaling 100 projects. To ensure each project provided sufficient data for robust analysis, we excluded projects with fewer than ten reported bugs. This criterion was crucial for enabling meaningful comparisons of bug localization techniques across projects with substantial debugging histories. After applying this filter, we retained 18 Frameworks, 29 Tools, and 42 Libraries, resulting in 89 projects for further analysis.
Filter 2: We identified overlaps where projects belonged to multiple categories (e.g., a tool functioning as a library). To resolve this, we meticulously reviewed official documentation and the functional roles of these projects, assigning each to a unique category. This refinement led to a final dataset of 86 projects and 1,795 bug reports, distributed as follows: 17 Frameworks (746 bugs), 28 Tools (455 bugs), and 41 Libraries (594 bugs).
To better understand the challenges in localizing deep learning bugs, we analyzed the performance of our baseline techniques across different categories of bugs in DL systems. We employed stratified random sampling to select a balanced number of bug reports from each category for performance comparison, ensuring fair representation and reducing selection bias. We repeated the process three times with different samples to ensure robustness. The results were averaged across these evaluations, providing reliable insights into the efficacy of the bug localization methods.